OpenStudy (anonymous):

6.Using complete sentences, describe the characteristics and parts of sigma notation. Please also provide a unique example.

OpenStudy (anonymous):

Lol. another question on sigma notation. Sigma notation is a series denoted by the Greek letter sigma with n=something, which is basically the number used to start. The top part of the sigma sign is where n goes to. For example, if n=1 and goes to infinity, then the sigma notation will look like this. \[\sum_{n=1}^{\infty}\] It will have a function or a series like \[\sum_{n=1}^{2}(1/n)=3/2\] Since (1/1)+(1/2)=(3/2). Hope this helps!

OpenStudy (anonymous):

it does! Helps ALOT! your a life saverrrr<3 thanks so much!!

OpenStudy (anonymous):

I have another question on sigma notation...if you would like to helpp? PLease! @berlingots

OpenStudy (unklerhaukus):

Sigma notation is used to denote a sum of finite terms The equation below the Sigma (n=-1) indicates the dummy variable (n), and its initial value (-1) The term above the Sigma indicates the maximum value of the dummy variable The terms to the right of the Sigma (1) may or may not be a function of n, and are to be evaluated over the range of the dummy variable, where the dummy variable increases in integers. \[\sum\limits_{n=-1}^{2}2^n=2^{-1}+2^0+2^1+2^2\] An example where the dummy variable does not occur in the terms to be evaluated \[\sum\limits_{n=1}^{4}1=1+1+1\]